Fermi National Laboratory
 

Accelerator Update for Jun 9, 2006 - June 12, 2006

 

Friday June 9

The day shift began with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Booster, Main Injector (MI), Recycler, Tevatron (TeV) conducting studies, and with NuMI and MiniBooNE taking beam.

Pbar experts changed their stacking rate to a three second cycle at 9:29 AM.

At 10:36 AM, the TeV quenched at sector B0. Operations discovered a B0 quadrupole (B0Q6) was the first magnet to quench. Investigations continued.

Operations prepared the Collider Detector at Fermilab (CDF) for access at 10:54 AM. D-Zero (D0) decided to use this as a quiet time.

CDF came out of access at 11:35 AM.

At 1:29 PM, the TeV turn on was delayed an hour while technicians replaced a shunt module for the sector A2 dump.

At 1:43 PM, the sector A2 cryo system cold compressor tripped off during a ramp test. An expert investigated and reported that the TeV had suffered another quench at A2.

Linac techs began work on the Klystron Debuncher water skid at 2:06 PM. They completed their work thirty minutes later. This work held off all downstream beam. Techs also worked on MV-1 and the H- Source and this held off beam for an another fifteen minutes.

Operations began preparing the Recycler to stash antiprotons at 5:03 PM.

Operations put the TeV into a dry squeeze at 5:03 PM.

The Recycler stash ended at 6:35 PM. There were five transfers. Pbar resumed stacking at 6:51 PM.

Operations and TeV experts put the TeV into shot setup at 7:43 PM.

Operations established store 4745 at 9:47 PM with an initial luminosity of 1.45E30. This was a 12x12 store (12 proton bunches by 12 antiproton bunches).

Saturday June 10

The midnight shift began with Operations monitoring store 4745 and a small stash, with Pbar stacking and conducting studies, with CDF and D0 using the colliding beam, Booster, MI, and the RR conducting studies, and with MiniBooNE and NuMI taking beam.

At 1:57 AM, a number of Recycler bulk power supplies tripped off. Operators turned them back on, but not before 0.2E10 of antiprotons were lost.

At 6:21 AM, the TeV suffered a quench at F sector. TeV experts were conducting studies with the remains of the store.

Operations put the TeV into a dry squeeze at 8:28 AM.

TeV experts began a study period at 9:58 AM.

Pbar resumed stacking at 2:54 PM.

Recycler began stashing at 5:25 PM. The stash ended and Pbar resumed stacking at 6:46 PM.

Operators prepared Pbar for an access at 7:35 PM to investigate a vacuum problem. Operations called in vacuum experts. They reported that tank 201 in the Debuncher had a number of leaks. They tried a vacuum seal, but it didn’t work. The tank vacuum will be allowed to come up to normal atmosphere. In the morning the experts will make repairs.

Sunday June 11

The midnight shift began with Operations monitoring a small stash, with Pbar off due to vacuum problems, with the TeV separators in conditioning mode, with Booster, MI, and RR conducting studies, and with MiniBooNE and NuMI taking beam.

TeV experts completed their studies at 3:19 AM.

At 6:18 AM, Operations reported that they were having problems keeping the NuMI beam permit up. Operators contacted an expert and later called in people to try to resolve a problem with the target cooling system.

At 6:56 AM, a MI-60 PLC fault took down the electrical safety system, which took everything down, including the Recycler. The stash was lost. Operations called in a safety system tech.

The safety system tech replaced a failed processor. Operators began a search and secure of MI at 9:43 AM.

Operators prepared Pbar for an access at 11:36 AM.

Operations established beam to MiniBooNE at 11:40 AM.

Operations prepared D0 for an access at 1:08 PM. Vacuum techs repaired the leaks on tank 201. It will take until tomorrow morning for the tank to complete its vacuum pump down.

NuMI experts began targets scans at 1:57 PM.

Operations established beam to NuMI at 2:05 PM.

Operations put the TeV into a dry squeeze at 3:45 PM.

At 5:20 PM, the Klystron Debuncher water skid tripped off for the first of two times in thirty-eight minutes. It reset both times.

At 9:12 PM, NuMI tripped off.

At 9:42 PM, Operations reported that no beam was getting to MI. Several power supplies had tripped off on a ground fault indication. Everything reset okay.

Operations reset the NuMI trip at 9:49 PM, but no beam appeared. Operations investigated and discovered beam valves had closed in Booster, which had suffered a vacuum burst.

At 9:52 PM, MI tripped off again on power supply ground faults (I:V519, I:H520, & I:V521). Everything reset. Operations contacted an expert.

At 11:01, MI tripped off again with I:H518 added to the list of ground faults. Everything reset. The expert told Operations to monitor the main bus current and that he would investigate further in the morning.

Monday June 112

The midnight shift began with Operations monitoring the stash, with Pbar off, with the TeV conducting separator conditioning and studies, with Booster, MI, and RR conducting studies, and with MiniBooNE and NuMI taking beam.

At 12:04 AM, NuMI tripped off due to Horn power supply rectifiers.

At 2:11 AM, the Klystron Debuncher water skid tripped off, was reset and then tripped again. It reset okay.

Operations reported at 2:14 AM, that the NuMI Horn rectifiers would not reset. Operators notified experts.

The Plans for Monday

The plans for today are to get NuMI back on and to, hopefully, establish a normal 36x36 store this evening.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/12/2006    email Fermilab